OpenStudy (anonymous):

find all the x values where relative maxima occur for the function g(x)=2x^3-3x^2-5 on the interval (-5,2) A. x = 1 B. x = 0 C. x = 0 and x = 1 D. None of the above.

OpenStudy (anonymous):

Differentiating the above equation..gives you the answer..!!

OpenStudy (anonymous):

Would be Like 6x^2-6x =0 and From here taking x common we Have,

OpenStudy (anonymous):

x=0 and 6x-6=0 or x=0 and x-1=0 or x=0 and x=1.

OpenStudy (anonymous):

Third Option :))

OpenStudy (anonymous):

thanks! :)